# Partnering with Generative AI: Experimental Evaluation of Human-Led and Model-Led Interaction in Human-AI Co-Creation

### v1

@@ Appendix C Examples of Ideas
Appendix D Correlations
Table 6. Means, standard deviations, and correlations
@@ Appendix D Correlations
  • Note: M and SD represent mean and standard deviation, respectively. * indicates p ¡ .05; ** indicates p ¡ .01.
Appendix E Rater Instructions: Idea Quality
Adapted from “An Empirical Investigation of the Impact of ChatGPT on Creativity” (Lee & Chung, adaptation).
@@ 3. General Guidelines
Appendix F Coding Scheme: Elaboration Depth
We based the coding on seven characteristics and coded whether these were added over the LLM input:
Characteristics for Idea Elaboration (CIEs)
  • Principle: User added a guiding rule/ordering that steers later choices.
  • Mechanism: User added how it works—a trigger, process, protocol, or algorithm.
  • Feature: User added a distinct user-facing capability or mode.
  • Constraint: User added a binding limit/requirement that gates behavior.
  • Audience/Context: User added the target users or setting where the idea applies.
  • Integration/Architecture (incremental adaptation): User integrated LLM input and added an incrementally new idea based on LLM input—combining, syncing, or restructuring components to extend/refine the same core concept.
  • Reframing/Pivot: User provided a radically new idea based on LLM input—replacing the core frame/value logic so the concept serves a different primary purpose, audience, or setting.
Elaboration Levels
For every interaction we coded these principles and applied a certain elaboration level based on them:
Level 1 — LLM-led echo (accepts LLM input as-is):
CIEs: 0 (pure acceptance/endorsement).
Trigger: Nothing new added.
Depth: None (endorsement).
Level 2 — Minor elaboration: Integrates details provided by LLM:
CIEs: $\geq$1 Feature or Mechanism, or a binding Constraint, Audience/Context; but already provided by AI, core frame unchanged.
Depth: Surface tweak.
Level 3 — Substantive elaboration (adds a shaping feature/mechanism):
Human CIEs: $\geq$1 Feature or Mechanism, or a binding Constraint, Audience/Context; but independently from AI.
Depth: Concrete conceptual detailing (feature + operational rule).
Level 4 — Integrative advancement (new idea):
Human CIE: Integration/Architecture (incremental adaptation).
User added an incrementally new idea based on LLM input—combining, syncing, or restructuring components to extend/refine the same core concept.
Depth move: Integration + Creation.
Level 5 — Human-led reframing (pivot + new organizing mechanism):
Human CIE: Reframing/Pivot.
User added a radically new idea based on LLM input—replacing the core frame/value logic so the concept serves a different primary purpose, audience, or setting.
Depth move: Conceptual pivot + radical innovation.
Appendix G Perceived Cognitive Workload
@@
Appendix H Within-Subject Analyses

### v2

@@ Appendix B Examples of Ideas (Initial Study)
Appendix C Correlations (Initial Study)
Table 8. Means, standard deviations, and of the main study
@@ Appendix C Correlations (Initial Study)
Note: M and SD represent mean and standard deviation, respectively. * indicates p <.05<.05; ** indicates p <.01<.01.
Appendix D Rater Instructions: Idea Quality
Adapted from Lee and Chung (2024).
@@ 3. General Guidelines
Appendix E Coding Scheme: Elaboration Depth
Table 9. Integrated Coding Scheme: Mapping Themes (Levels) to Codes (CIEs)
Theme (Level) Applicable Codes (CIEs) Description & Criteria Depth Move
Level 1
LLM-led Echo
None (0 CIEs) Pure Acceptance. User accepts LLM input as-is. No new characteristics are added. Endorsement
Level 2
Minor Elaboration
Feature
Mechanism
Constraint
Audience/Context
Integration of AI details. User adds codes ($\geq$1), but these details were already provided by the AI. Core frame remains unchanged. Surface Tweak
Level 3
Substantive Elaboration
Feature
Mechanism
Constraint
Audience/Context
Principle
Independent Detailing. User independently adds codes ($\geq$1) that were not present in the AI input. Adds concrete conceptual detailing. Conceptual Detailing
Level 4
Integrative Advancement
Integration / Architecture Incremental New Idea. User combines, syncs, or restructures components to extend the same core concept based on LLM input. Integration + Creation
Level 5
Human-led Reframing
Reframing / Pivot Radical Innovation. User replaces the core frame or value logic so the concept serves a different primary purpose or audience. Conceptual Pivot
Note. The coding process followed a two-step logic: First, we analyzed every user interaction to identify the presence of any of the seven codes (CIEs), such as adding a Feature, Mechanism, or Constraint. Second, we assigned a thematic Level (1–5) based on the assigned codes and their source (AI vs. Human).
Appendix F Perceived Cognitive Workload (Initial Study)
@@
Appendix G Within-Subject Analyses (Initial Study)
@@
Appendix H Robustness Checks
We assessed the assumptions for all regression models across both studies, examining posterior predictive fit, homogeneity of variance, influential observations, and normality of residuals (including random effects for the linear mixed models). While assumptions were largely satisfied, some models showed some evidence of potential residual non-normality or heteroscedasticity. Detailed diagnostic plots for all dependent variables can be found in our GitHub. To ensure robustness, we supplemented the parametric analyses with non-parametric alternatives: Kruskal–Wallis tests with Dunn’s post hoc comparisons for the linear regression models, and ordinal mixed models for the linear mixed models. In the Dunn’s post hoc analyses, we primarily report unadjusted p-values. Additionally, we report holm-adjusted p-values that were corrected across all pairwise comparisons simultaneously (even the comparisons we do not report). Because non-parametric procedures do not allow contrast specification, this yields particularly conservative p-values. Additionally, we applied a logit transformation to the cosine similarity values, which are bounded between 0 and 1, to better satisfy distributional assumptions. Importantly, results from the robustness checks were consistent with those from the parametric hypothesis tests and confirm our main conclusions.
H.1. Study 1
Table 14. Robustness check: Hypothesis tests for idea quality with non-parametric Kruskal–Wallis test
Hyp. Comparison $z$ $r$ $p$ $p_{\text{adj}}$
Idea quality
H1a Q Question-mode >> C Control $2.23$ $0.16$ $013^{*} .$090.090
H1b S Suggestion-mode >> C Control $1.07$ $0.08$ .143.143 .500.500
H1c M Model-led >> C Control $3.88$ $0.28$ $^{***} .001.001^{**}$
H1d V Vanilla >> C Control $0.21$ $0.02$ .415.415 .500.500
Notes: An omnibus Kruskal–Wallis test revealed significant differences in idea quality across interaction modes ($\chi^{2}(4)=20.50$, $p<.001$). Pairwise comparisons were conducted using Dunn’s post hoc test. Reported are $z$-statistics, Holm-adjusted one-sided $p$-values (across all 10 pairwise comparisons), and $r$ effect sizes; only hypothesis-relevant comparisons are shown. Significance levels: * $p<.05$, ** $p<.01$, *** $p<.001$. The non-significant adjusted $p$-value for idea quality is likely a consequence of correcting for all 10 pairwise comparisons. Only hypothesis-relevant comparisons are reported.
Table 15. Robustness check: Hypothesis tests for idea diversity with non-parametric Kruskal–Wallis test
Parametric test (logit transformed) Dunn’s test
Hyp. Comparison $t$ (df) $p$ $d$ $z$ $p$ $p_{\text{adj}}$ $r$
Idea diversity
H2a Q Question-mode >> M Model-led 5.20(481)5.20\;(481) $001^{***}$ $0.74$ $5.72$ $^{***} <.001<.001^{***}$ $0.41$
H2b Q Question-mode >> V Vanilla 0.27(481)0.27\;(481) .791.791 $0.04$ $0.14$ .444.444 .500.500 $0.01$
H3a S Suggestion-mode >> M Model-led 1.45(481)1.45\;(481) .223.223 $0.21$ $2.40$ $008^{**} .$$024^{*}$ $0.17$
H3b S Suggestion-mode >> V Vanilla 3.42(481)-3.42\;(481) .999.999 0.49-0.49 3.11-3.11 .999.999 .993.993 0.23-0.23
Notes: An omnibus Kruskal–Wallis test revealed significant differences in idea diversity across interaction modes ($\chi^{2}(4)=46.11$, $p<.001$). Pairwise comparisons were conducted using Dunn’s post hoc test. Reported are $z$-statistics, unadjusted and Holm-adjusted one-sided $p$-values (across all 10 pairwise comparisons), and $r$ effect sizes; only hypothesis-relevant comparisons are shown. Significance levels: * $p<.05$, ** $p<.01$, *** $p<.001$.
Table 16. Robustness check: Hypothesis tests for perceived idea ownership with non-parametric Kruskal–Wallis test
Hyp. Comparison $z$ $p$ $p_{\text{adj}}$ $r$
Perceived idea ownership
H4a Q Question-mode >> M Model-led $3.88$ $^{***} <.001<.001^{***}$ $0.28$
H4b Q Question-mode >> V Vanilla $3.42$ $01^{***} .00$$02^{**}$ $0.25$
H5a S Suggestion-mode >> M Model-led $0.93$ .177.177 .500.500 $0.07$
H5b S Suggestion-mode >> V Vanilla $0.52$ .301.301 .500.500 $0.04$
Notes: An omnibus Kruskal–Wallis test revealed significant differences in perceived idea ownership across interaction modes ($\chi^{2}(4)=61.96$, $p<.001$). Pairwise comparisons were conducted using Dunn’s post hoc test. Reported are $z$-statistics, unadjusted and Holm-adjusted one-sided $p$-values (across all 10 pairwise comparisons), and $r$ effect sizes; only hypothesis-relevant comparisons are shown. Significance levels: * $p<.05$, ** $p<.01$, *** $p<.001$.
H.2. Study 2
Table 17. Robustness check: Hypothesis tests for idea quality with ordinal mixed model
Hyp. Comparison Estimate SE $z$ $p$
Idea quality
H1a Q Question-mode >> C Control $0.86$ $0.16$ $5.55$ $001^{***}$
H1b S Suggestion-mode >> C Control $0.02$ $0.15$ $0.13$ .899.899
H1c M M-led (1-shot) >> C Control $0.87$ $0.14$ $6.21$ $001^{***}$
H1d M-I M-led (iter.) >> C Control $0.68$ $0.15$ $4.38$ $001^{***}$
H1e V Vanilla >> C Control 0.17-0.17 $0.16$ 1.09-1.09 .899.899
Notes: Pairwise comparisons of interaction modes on idea quality using an ordinal mixed model approach (cumulative link mixed model with random intercepts for idea and evaluator). Reported are estimates on the log-odds scale, standard errors, $z$-statistics, and Holm-adjusted $p$-values (one-sided, right-tailed tests). Estimates represent the change in log-odds of being in a higher quality category compared to the control condition. The overall support for the hypotheses is robust and equivalent to the parametric linear mixed model test. Significance levels: * $p<.05$, ** $p<.01$, *** $p<.001$.
Table 18. Robustness check: Hypothesis tests for idea diversity with logit-transformed values
Parametric test (logit transformed) Dunn’s test
Hyp. Comparison $t$ (df) $p$ $d$ $z$ $p$ $p_{\text{adj}}$ $r$
Idea diversity
H2a Q Question-mode >> M M-led (1-shot) 3.34(634)3.34\;(634) $002^{**}$ $0.44$ $3.28$ $01^{**} .00$$04^{**}$ $0.21$
H2b Q Question-mode >> M-I M-led (iter.) 1.11(634)1.11\;(634) .396.396 $0.16$ $0.95$ .171.171 .218.218 $0.07$
H2c Q Question-mode >> V Vanilla 3.22(634)-3.22\;(634) .999.999 0.47-0.47 2.51-2.51 .994.994 .969.969 0.18-0.18
H3a S Suggestion-mode >> M M-led (1-shot) 8.20(634)8.20\;(634) $001^{***}$ $1.06$ $7.72$ $^{***} <.001<.001^{***}$ $0.49$
H3b S Suggestion-mode >> M-I M-led (iter.) 5.51(634)5.51\;(634) $001^{***}$ $0.78$ $4.97$ $^{***} <.001<.001^{***}$ $0.35$
H3c S Suggestion-mode >> V Vanilla 1.12(634)1.12\;(634) .396.396 $0.16$ $1.46$ .073.073 .218.218 $0.10$
Notes: An omnibus Kruskal–Wallis test revealed significant differences in idea diversity across interaction modes ($\chi^{2}(5)=117.72$, $p<.001$). Pairwise comparisons were conducted using parametric (linear mixed model) and non-parametric (Dunn’s test) approaches. For the parametric test, reported are $t$-statistics with degrees of freedom, Holm-adjusted one-sided $p$-values (across the 6 contrasts), and Cohen’s $d$ effect sizes. For the non-parametric test, reported are $z$-statistics, Holm-adjusted one-sided $p$-values (across all 15 pairwise comparisons), and $r$ effect sizes; only hypothesis-relevant comparisons are shown. Significance levels: * $p<.05$, ** $p<.01$, *** $p<.001$.
Table 19. Robustness check: Hypothesis tests for perceived idea ownership with non-parametric Kruskal–Wallis test
Hyp. Comparison $z$ $p$ $p_{\text{adj}}$ $r$
Perceived idea ownership
H3a Q Question-mode >> M M-led (1-shot) $7.82$ $^{***} <.001<.001^{***}$ $0.56$
H3b Q Question-mode >> M-I M-led (iter.) $3.52$ $^{***} .001.001^{**}$ $0.22$
H3c Q Question-mode >> V Vanilla $5.57$ $^{***} <.001<.001^{***}$ $0.40$
H4a S Suggestion-mode >> M M-led (1-shot) $5.93$ $^{***} <.001<.001^{***}$ $0.42$
H4b S Suggestion-mode >> M-I M-led (iter.) $1.43$ .076.076 .076.076 $0.09$
H4c S Suggestion-mode >> V Vanilla $3.70$ $^{***} .001.001^{**}$ $0.27$
Notes: An omnibus Kruskal–Wallis test revealed significant differences in perceived idea ownership across interaction modes ($\chi^{2}(5)=130.38$, $p<.001$). Pairwise comparisons were conducted using Dunn’s post hoc test. Reported are $z$-statistics, Holm-adjusted one-sided $p$-values (across all 15 pairwise comparisons), and $r$ effect sizes; only hypothesis-relevant comparisons are shown. Significance levels: * $p<.05$, ** $p<.01$, *** $p<.001$.
Appendix I Analyses Validation Study
Purchase intent means with confidence intervals by condition Point-and-interval plot of purchase intent mean ratings by condition, with vertical error bars showing ninety-five percent confidence intervals. Conditions on the horizontal axis are question-mode, suggestion-mode, model-led, generation-mode, vanilla, and control. Question-mode and model-led have the highest mean purchase intent values, suggestion-mode has the lowest mean, and generation-mode, vanilla, and control lie in between.
Bar chart showing purchase intent means with 95% confidence intervals across five conditions. The y-axis ranges from 2.1 to 2.6. Question-Mode shows the highest mean (approximately 2.55), followed by Suggestion-Mode (approximately 2.5), Model-Led Generation-Mode (approximately 2.45), Vanilla (approximately 2.35), and Control (approximately 2.3). Error bars indicate 95% confidence intervals for each condition, with some overlap between adjacent conditions.
Figure 8. Purchase intent varies across AI interaction modes. Mean purchase intent ratings with 95% confidence intervals across five conditions (scale: 1–7).
Cognitive workload bars and perceived attribution of creative thinking line Combined bar and line chart by condition (question-mode, suggestion-mode, model-led, generation-mode, vanilla, control). Bars show mean cognitive workload on a zero to one hundred scale with error bars. A line with markers shows attribution of creative thinking on a zero to one hundred scale where zero means human did most of the thinking and one hundred means artificial intelligence did most of the thinking. Attribution is highest in generation-mode and lower in question-mode and control. Cognitive workload is higher in the human-led modes and control than in the model-led conditions.
Figure 9. Perceived cognitive workload (bars) and attribution of creative thinking (blue line) across conditions. Cognitive workload was measured using NASA-TLX on a 0–100 scale. Attribution of creative thinking was assessed with the question “Who did more of the creative thinking required to develop the idea?” (0 = Human did most of the thinking, 100 = AI did most of the thinking). Error bars represent standard errors. The perceived cognitive workload showed similar patterns to the initial study, with human-led conditions (question-mode $M=57.14$, $SD=17.91$; suggestion-mode $M=52.42$, $SD=21.87$) showing higher values compared to the model-led conditions (model-led iterative $M=46.01$, $SD=23.03$; model-led one-shot $M=44.04$, $SD=22.46$), aligning with perceived contribution (human vs. AI) except for the model-led one-shot condition ($M=39.98$, $SD=25.72$) which showed similar perceived AI contribution to suggestion-mode ($M=40.34$, $SD=27.12$).